Shift report 20020118 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 18 2002 - 08:18:55 EST

  • Next message: Betty McBreen: "Inbound ssh to BNL's computers"

    ------------------------------------------------------------
    Date: Fri Jan 18 08:18:55 2002, Shift: 00:00-08:00
    Supervisor: JH Lee, with: 
    Log book#12 pages 77 - 81
    
    Shift Summary:
    ------------------------------------------------------------
    1. Physics Beam since 02:50 
       started at rates of
       - B/Y = 25/40*10^11
       - In-el 18K/sec
       - 15 trigger3/sec 
       
    2. Data:
       - MRS at 90 6B
       - FS at 3 1/5B
       - run 6648: 47K trigger3 210K trigger6 in 100 min.
         (scale down trigger3:trigger6 = 1:5)   
         run 6656: 60K trigger3 and still running
       
    
    3. Problems:
    
       T2 readout problem:  
       - Couldn't take data with T2 in DAQ after run 6648.
       - Recycling power of T2 readout board couldn't fix 
         the problem.  
       - Now running with only MRS (run 6656) 
       - It needs to be investigated after the current store.
    
       D3 Trip and reset.
                              
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Jan 18 2002 - 08:19:25 EST